OpenStudy (anonymous):

Solve this system of equations 5x+4y=-14 3x+6y=6

OpenStudy (e.mccormick):

Do you ned to use any specific method?

OpenStudy (squirrels):

Okay for this equation we can either use substitution or elimination method. I think elimination is easier though.

OpenStudy (squirrels):

@e.mccormick plx

OpenStudy (e.mccormick):

Well, for this one, substitution might be easier... solving for x in the second one. \(\ddot \smile\) But both work really well.

OpenStudy (squirrels):

Oh yes that's correct, my mistake. What he said^^^

OpenStudy (squirrels):

But yes, is there a specific method they want you to use? @Keridavis

OpenStudy (texaschic101):

I agree....divide each term of the 2nd equation by 3. Then find what you would sub in for x.

OpenStudy (e.mccormick):

Keri, Keri, blueberry, where have your replies gone? So many here to help you, but from you not a song.

OpenStudy (the_fizicx99):

^

OpenStudy (anonymous):

? I need something simple and easy to work the problem

OpenStudy (e.mccormick):

OK. Then what do you get if you solve the second equation for x?

OpenStudy (anonymous):

Thats the problem, i have no idea how.....

OpenStudy (e.mccormick):

2a + 4b = 8 Solve for a means use math to get a alone. 2a + 4b - 4b = 8 - 4b 2a = 8 - 4b 2a/2 = (8 - 4b)/2 a = 4 - 2b You need to do the same sort of thing.

OpenStudy (e.mccormick):

Once you have what x is in terms of y, or sole for x, you can use that answer to replace the x in the first equation. That will lead to knowing what y is. Then you can use y to find x. That process is called substitution. In this case, it is easier.

OpenStudy (anonymous):

oh. I kinda get it.

OpenStudy (e.mccormick):

So, if you solve the second equation for x, what do you get?

OpenStudy (anonymous):

Gimme a minute xD

OpenStudy (e.mccormick):

One, one thousand Two, one thousand Three, one thousand Four, one thousand Gonna take me more than a minute to type 60 seconds. =P

OpenStudy (anonymous):

xD Sorry, And I got x= 3-3y *hopes I'm right*

OpenStudy (e.mccormick):

Yes. So, that is a value of x. Not a number, but still 100% valid. So replace x in the first equation with (3-3y) and what do you get?

OpenStudy (anonymous):

Wait, So it will be,.. " 5(3-3y)+4y = -14???

OpenStudy (e.mccormick):

Exactly! Once you solve that, you will know what y is.

OpenStudy (anonymous):

OMG! Thank you so much! No one could explain this to me at alll!

OpenStudy (e.mccormick):

And once you know exactly what y is, you put that into either equation and find x.

OpenStudy (e.mccormick):

An easy way to check the answer is to try y in both and make sure you get the same x. Another is to graph both lines and they should meet at the (x,y) point that solves this.

OpenStudy (anonymous):

oh god.. I got a decimal...

OpenStudy (e.mccormick):

Hmmm.... lets see. 5(3-3y)+4y = -14 15-15y+4y = -14 15-11y = -14 -11y = -14-15 -11y = -29 Well, you can leave it as a fraction. y=29/11

OpenStudy (e.mccormick):

Hmmm... but when I graph it... odd. Let me see if we made a mistake somwhwere.

OpenStudy (e.mccormick):

3x+6y=6 3x=6-6y x=2-2y Oops. Did not catch that earlier.

OpenStudy (anonymous):

o.O? Oh.. I'm lost..

OpenStudy (e.mccormick):

Try it again with (2-2y) rather than (3-3y). Just a small mistake.

OpenStudy (anonymous):

Wait.. Where did I make that mistake??

OpenStudy (e.mccormick):

it is not x = 3 - 3y Here is the full math of it. Check yours: 3x+6y=6 3x+6y-6y =6-6y 3x=6-6y 3x/3=(6-6y)/3 x=6/3-6y/3 x=2-2y

OpenStudy (e.mccormick):

It is probably in this step: x=6/3-6y/3 You went to 3 instead of 2. A minor logical error. We all make them at times.

OpenStudy (anonymous):

Oh, I see now. thnk you.